OpenStudy (anonymous):

HELPPPPPPPPPPP! PLEASEEEE If the length of each side of a cube is doubled, is the surface area doubled? Use an example to support your answer.

OpenStudy (ash2326):

@Kayaya Do you know the formula relating surface area of a cube to its side?

OpenStudy (anonymous):

yes

OpenStudy (kaederfds):

nope. Example: length=3 surface area= 6.3^2 = 6.9= 54 length=6 surface area= 6.6^2= 6.36= 216

OpenStudy (anonymous):

thx